# Probing Spin Defects via Single Spin Relaxometry

### v3

@@ Abstract
Spin defects in solids offer promising platforms for quantum sensing and memory due to their long coherence times and optical addressability. Here, we integrate a single nitrogen-vacancy (NV) center in diamond with scanning probe microscopy to discover, read out, and spatially map arbitrary spin-based quantum sensors at the nanoscale. Using the boron vacancy (V) center in hexagonal boron nitride—an emerging two-dimensional spin system—as a model, we detect its electron spin resonance indirectly via changes in the spin relaxation time () of a nearby NV center, eliminating the need for optical excitation or fluorescence detection of the V. Cross-relaxation between NV and V ensembles significantly reduces NV , enabling quantitative nanoscale mapping of defect densities beyond the optical diffraction limit and clear resolution of hyperfine splitting in isotopically enriched . Our method demonstrates interactions between 3D and 2D spin sensors, establishing NV centers as versatile probes for characterizing otherwise inaccessible spin defects.
@@ Introduction
Currently, hexagonal boron nitride (hBN) is the only known two-dimensional material hosting optically active spin defects, making it uniquely promising for surface-proximal quantum sensing [16]. In addition to hBN with naturally occurring isotope ratios (hBN), isotopically purified allows clear observation of the hyperfine structure, giving insight into electron-nuclear interactions as well as an increase in magnetic field sensitivity [8, 19]. Nevertheless, a major limitation remains: spin-active quantum defects in two-dimensional materials that emit at telecom wavelengths—a critical requirement for integration into fiber-based quantum systems—have yet to be identified [24]. Overcoming this barrier necessitates the development of a scalable, efficient method to discover and read out new spin-active quantum defects, particularly those operating in the near-infrared (NIR)/telecom spectral range, where single-photon detection remains costly and technically challenging.
In this work, we use the boron vacancy () center in hBN and as a model system to demonstrate that NV-based relaxometry can read out and spatially map arbitrary spin-active quantum defects. The center is a point defect consisting of a missing boron atom in the hBN lattice, exhibiting a spin-1 ground state with optically detected magnetic resonance (ODMR) at room temperature [22, 21, 9, 19]. By sweeping the magnetic field through the NV- spin cross-relaxation (CR) condition and monitoring the resulting changes in the NV center’s spin relaxation time, we indirectly detect the boron vacancy’s electron spin resonance (ESR) spectrum in hBN—eliminating the need for direct optical excitation, microwave driving, or emission-based readout of defects. In addition, we demonstrate the ability of cross-relaxation spectroscopy to resolve the hyperfine structure in , revealing nuclear spin information. Finally, we perform nanoscale spatial mapping of the cross-relaxation signal, revealing spatial heterogeneity in spin noise and quantitatively correlating this noise with V defect densities. Critically, our method addresses a key limitation of conventional defect-mapping techniques (e.g., Raman, electron microscopy) that cannot distinguish neutral from negatively charged boron vacancies, demonstrating that fewer than 10% of the vacancies are negatively charged and thus suitable as quantum sensors. These findings emphasize the versatility of scanning NV corss-relaxometry as a powerful method for probing emerging quantum spin systems, especially those with limited optical accessibility.
@@ Results
The total relaxation rate of the NV center can here be modeled as a contribution from three terms:
@@ Discussion
Materials and Methods
@@
The crystal flakes were grown by the atmospheric pressure high temperature (APHT) method, as previously described in detail [32, 26]. Briefly, the process starts by mixing high purity elemental boron, enriched in the 10B isotope to 99%, with nickel and chromium with weight ratios of 4:48:48, respectively. The mixture is then heated under a blanket of nitrogen enriched in the 15N isotope to 99.5% at a pressure of 850 torr to 1550°C, to produce a homogeneous molten solution. After 24 hours, the solution is slowly cooled at 4°C/h, to 1500°C, then at 50°C/h to 1300°C, and 200°C/h to room temperature. The solubility is decreased as the temperature is reduced, causing crystals to precipitate on the surface of the metal. The flakes were exfoliated from the metal with thermal release tape. The free-standing crystalline flakes were typically 10 to 20 m thick. To create boron vacancies, the flakes were neutron-irradiated. Samples were irradiated in The Ohio State University nuclear reactor, operated at a power 300 kW for 1 hour, corresponding to neutron fluences of neutron/cm2. Prior to the cross-relaxometry experiments, micrometer-scale flakes were exfoliated from irradiated macroscopic crystals using standard mechanical exfoliation. This process yielded flakes with thicknesses of approximately 50 nm, used for the measurements shown in Fig. 2, and approximately 15 nm, used for the data presented in Fig. S26.
@@
All data needed to evaluate the conclusions of this paper are available in the Article or the Supplementary Information. The experimental and simulation data generated in this study are provided in the Source Data file. All raw data relevant to this work are available from the corresponding author upon reasonable request. Source data are provided with this paper.
@@ Acknowledgments:
The scanning NV microscopy, hBN synthesis, and nanofabrication were supported by the Center for Nanophase Materials Sciences, (CNMS), which is a US Department of Energy, Office of Science User Facility at Oak Ridge National Laboratory. Spin relaxation measurements were supported by the Laboratory Directed Research and Development Program of Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the U.S. Department of Energy. The RF controls were support by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. Support for crystal growth was provided by the Office of Naval Research, award number N00014-22-1-2582. Neutron irradiation of the crystals was supported by the U.S. Department of Energy, Office of Nuclear Energy under DOE Idaho Operations Office Contract DE-AC07-051D13417 as part of a Nuclear Science User Facilities experiment. This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).
@@
Materials & Correspondence:
Correspondence and requests for materials should be addressed to Huan Zhao

### v4

# Probing Boron Vacancy Defects in hBN via Single Spin Relaxometry

@@ Abstract
Spin defects in solids offer promising platforms for quantum sensing and memory due to their long coherence times and optical addressability. Here, we integrate a single nitrogen-vacancy (NV) center in diamond with scanning probe microscopy to detect, read out, and spatially map spin-based quantum sensors at the nanoscale. Using the boron vacancy (V) center in hexagonal boron nitride—an emerging two-dimensional spin system—as a model, we detect its electron spin resonance indirectly via changes in the spin relaxation time () of a nearby NV center, eliminating the need for optical excitation or fluorescence detection of the V. Cross-relaxation between NV and V ensembles significantly reduces NV , enabling quantitative nanoscale mapping of defect densities beyond the optical diffraction limit and clear resolution of hyperfine splitting in isotopically enriched . Our method demonstrates interactions between spin sensors in 3D and 2D materials, establishing NV centers as versatile probes for characterizing otherwise inaccessible spin defects.
@@ Introduction
Currently, hexagonal boron nitride (hBN) is the only known two-dimensional material hosting optically active spin defects, making it uniquely promising for surface-proximal quantum sensing [16, 25]. In addition to hBN with naturally occurring isotope ratios (hBN), isotopically purified allows clear observation of the hyperfine structure, giving insight into electron-nuclear interactions as well as an increase in magnetic field sensitivity [8, 19]. More broadly, the discovery and characterization of new spin-active defects in low-dimensional materials—some of which may ultimately emit in technologically relevant spectral ranges such as the near-infrared or telecom—requires experimental tools that do not rely on defect-specific optical readout. A key outstanding challenge is therefore the ability to directly image the spatial distribution, density, and charge state of spin-active defects at the nanoscale. Conventional optical and structural probes are typically diffraction-limited, thickness-averaged, or unable to distinguish spin-active from spin-inactive charge states, masking nanoscale inhomogeneities that govern decoherence, spin–spin interactions, and device functionality. These limitations motivate the development of a nanoscale, charge-state-selective spin-density imaging technique that is independent of the optical properties of the target spin.
In this work, we use the boron vacancy () center in hBN and as a model system to demonstrate that NV-based relaxometry can read out and spatially map arbitrary spin-active quantum defects. The center is a point defect consisting of a missing boron atom in the hBN lattice, exhibiting a spin-1 ground state with optically detected magnetic resonance (ODMR) at room temperature [23, 22, 9, 19]. By sweeping the magnetic field through the NV- spin cross-relaxation (CR) condition and monitoring the resulting changes in the NV center’s spin relaxation time, we indirectly detect the boron vacancy’s electron spin resonance (ESR) spectrum in hBN—eliminating the need for direct optical excitation, microwave driving, or emission-based readout of defects. In addition, we demonstrate the ability of cross-relaxation spectroscopy to resolve the hyperfine structure in , revealing nuclear spin information. Finally, we perform nanoscale spatial mapping of the cross-relaxation signal, revealing spatial heterogeneity in spin noise and quantitatively correlating this noise with V defect densities. Critically, our method addresses a key limitation of conventional defect-mapping techniques (e.g., Raman, electron microscopy) that cannot distinguish neutral from negatively charged boron vacancies, demonstrating that fewer than 10% of the vacancies are negatively charged and thus suitable as quantum sensors. These findings emphasize the versatility of scanning NV cross-relaxometry as a powerful method for probing emerging quantum spin systems, especially those with limited optical accessibility.
@@ Results
The total relaxation rate of the NV center can here be modeled as a contribution from three terms [21]:
@@ Discussion
During revision of this manuscript, we became aware of a related study published by H. Sun et al. [45] reporting nanoscale imaging of spin defects in two-dimensional materials. Their results are complementary to those reported here.
Methods
@@
The crystal flakes were grown by the atmospheric pressure high temperature (APHT) method [33, 27]. Briefly, the process starts by mixing high purity elemental boron, enriched in the 10B isotope to 99%, with nickel and chromium with weight ratios of 4:48:48, respectively. The mixture is then heated under a blanket of nitrogen enriched in the 15N isotope to 99.5% at a pressure of 850 torr to 1550°C, to produce a homogeneous molten solution. After 24 hours, the solution is slowly cooled at 4°C/h, to 1500°C, then at 50°C/h to 1300°C, and 200°C/h to room temperature. The solubility is decreased as the temperature is reduced, causing crystals to precipitate on the surface of the metal. The flakes were exfoliated from the metal with thermal release tape. The free-standing crystalline flakes were typically 10 to 20 m thick. To create boron vacancies, the flakes were neutron-irradiated. Samples were irradiated in The Ohio State University nuclear reactor, operated at a power 300 kW for 1 hour, corresponding to neutron fluences of neutron/cm2. Prior to the cross-relaxometry experiments, micrometer-scale flakes were exfoliated from irradiated macroscopic crystals using standard mechanical exfoliation. This process yielded flakes with thicknesses of approximately 50 nm, used for the measurements shown in Fig. 2, and approximately 15 nm, used for the data presented in Fig. S26.
@@
The raw data of the main text figures are available in the Zenodo database [36]. Further data are available from the corresponding author upon request.
@@ Acknowledgments:
The scanning NV microscopy, hBN synthesis, and nanofabrication were supported by the Center for Nanophase Materials Sciences, (CNMS), which is a US Department of Energy, Office of Science User Facility at Oak Ridge National Laboratory. Spin relaxation measurements were supported by the Laboratory Directed Research and Development Program of Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the U.S. Department of Energy. The RF controls were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. Support for crystal growth was provided by the Office of Naval Research, award number N00014-22-1-2582. Neutron irradiation of the crystals was supported by the U.S. Department of Energy, Office of Nuclear Energy under DOE Idaho Operations Office Contract DE-AC07-051D13417 as part of a Nuclear Science User Facilities experiment. This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).
@@
Figure Legends
Fig. 1: NV relaxometry at CR condition with in hBN: (a) A scanning NV microscope containing a single NV center is brought near a 90-300 nm thick hBN sample that has been irradiated with He ions to create an ensemble of centers. The NV center’s photoluminescence (PL) is collected using a confocal optical setup. A microwave field is supplied to both the NV and the nearest centers by an antenna positioned near the cantilever tip. When the microwave frequency matches the spin resonance frequency of the NV or centers, transitions between the ground state sublevels states are driven coherently. (b) Calculated spin resonance dispersions of the NV and centers with the magnetic field oriented 29° from the surface normal. Cross-relaxation occurs near 127 G bias, where the NV transition overlaps with the transition, highlighted by the green dot. (c) CW-ODMR spectra of the NV center before (green curve) and after (purple curve) engaging with the hBN sample, and of the centers in hBN (red curve), measured under the cross-relaxation condition shown in (b). All spectra are fitted using Lorentzian functions. The left and right vertical axes correspond to the NV and data, respectively. Only the transition of the NV is shown. The arrows denote via their colors which vertical axis each trace is plotted against. (d) NV spin measurement in non-CR condition (purple curve, fitted ms) and CR condition (red curve, s), both measured after engaging the hBN sample. Curves are single exponential fits .
Fig. 2: -MR detection of hyperfine structure: (a) CW-ODMR spectrum of centers in showing four hyperfine-split lines, fitted with Lorentzian functions (red). A magnetic field of 123 G is applied out-of-plane. (b) Extracted ODMR center frequencies of the NV center (red) and centers (black) as a function of magnetic field, with solid lines representing fitted curves. The hyperfine line positions (dashed) are calculated using the average , revealing four expected cross-relaxation conditions corresponding to the crossover points between the NV and transitions. (c) NV single- -MR detection of hyperfine structure by sweeping the magnetic field through the four CR conditions. The magnetic field is converted to frequency detuning between the NV and transitions and plotted in units of in (67 MHz). The data are fitted using a sum of four Lorentzian functions, where dips 1 and 3 share one set of parameters, and dips 2 and 4 share another (fit shown in red). Here, the free evolution time is chosen to be 2 ms, given that the tip’s non-CR measured immediately before this -MR measurement is ms (see Supplementary Figure S23a). (d) NV single- -MR measurement of hBN under varying magnetic field. The field axis is converted to frequency detuning and plotted in units of in hBN (44 MHz). Data are fitted using a Lorentzian function (red curve). The in this case is 0.7 ms, and the tip’s measured at the non-CR condition is ms (see Supplementary Figure S23b). Error bars of (c) and (d) represent normalized shot noise level. (e) A short- NV tip’s spin relaxation curve before engaging the sample, showing of s. (f) NV values measured as a function of magnetic field (blue symbols with error bars), showing four distinct cross-relaxation dips. The curve is fitted with a sum of Lorentzian functions and plotted in units of in (67 MHz). Error bars represent uncertainties of fitted values.
Fig. 3: Nanoscale imaging of density in hBN using NV- cross relaxometry: (a) Optical microscope image of hBN sample on gold surface, with different colors corresponding to variations in the hBN thickness. The light-colored area in the center has thickness 90 nm compared to the surrounding pink/purple areas with thickness 190 nm. Here a portion of the right edge of the light-colored area is highlighted in a blue polygon which was measured by the scanning NV in panel (f). (b) Raman spectroscopy map showing the ratio of the peak to the peak across the sample area displayed in panel (a). (c) SRIM simulations of the ion energy losses (normalized per ion) of 30 keV He+ (10,000 ions simulated) for 90 nm and (d) 250 nm hBN on gold, suggesting thickness-dependent non-uniform defect density distribution. (e) Monte Carlo simulation of the cross-relaxometry contribution to the NV relaxation rate as a function of the density and three different NV-to-sample distances. Horizontal gray bar shows values of associated with range of the color scale in (f). (f) density map, obtained as a single- spatial scan over a free-evolution time s which was converted to density via the Monte Carlo simulation in panel (e). The size of each pixel is 100 nm. The measurement was conducted at the CR condition. Inset: The profile of density as a function of position, over the line indicated by the white dashed line in (f). The data is fit to the edge-spread function (red) yielding a step width of 46 nm.